User Tools

Site Tools


handover:r1724

R1724

Hobart

Module: HOB+0075, 4745.416 GB

  • 17:00 UT - Before observation checklist delays were stated at the wrong time, should have been 16:45UT (Tiege)
  • 17:00 UT - Experiment started OK (Tiege)
  • 1630UT Log missing ~20min of data due to fs hdd at hobart being full - fixed (JS)

Katherine

Module: GSFC+022, 3095.901 GB

  • 17:00 UT - Before observation checklist delays were stated at the wrong time, should have been 16:45UT (Tiege)
  • 17:00 UT - Experiment started OK (Tiege)
  • 17:01 UT - Delay difference shifted from ~0.3 microseconds to ~0.619 however it seems stable at this delay difference. (Tiege)
  • 17:25 UT - Delay continued to drift, halted schedule, Reset the DBBC, restarted observations, missed scans 026-1715b & 026-1721, prior scans likely unreliable. (Tiege)
  • 17:39 UT - Drifting happened again, not stabilising after 10+mins, halt schedule and rebooted DBBC, reconfigured and restarted schedule. Missed scans 026-1735b and 026-1738 (Tiege)
  • 18:07 UT - Yet again drifting occuring badly, tried doing a hard reboot of the DBBC using the ipswitch, currently monitoring, appears to be still drifting. missed 026-1808b through to 026-1822a (Tiege)
  • 18:43 UT - Called Warren (on-call) as delay was still consistently drifting, setup Kathernine again from scratch, reset field system, mk5, dbbc etc. restarted schedule at 19:00 UT, still appears to be drifting, Warren directed me to leave it until the morning. Scans missed 026-1843c through to 026-1854b (Tiege)
  • 2210-2229 UT - No data recorded while adjusting maser settings. Modulation turned off & no further clock jumps/slips noted. (JMc)

Yarragadee

Module: USN-0049, 0.000 GB

  • 17:00 UT - Before observation checklist delays were stated at the wrong time, should have been 16:45UT (Tiege)
  • 17:00 UT - Experiment started OK (Tiege)
  • 06:10 UT - Power outage at Yarragadee, requiring restart of computers on site. Missed scans 027-0606 through 027-0638b (Ross)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r1724.txt · Last modified: 2016/01/27 16:46 by Tiege McCarthy